# Date of publication xxxx 00, $0$, date of current version xxxx 00, $0$. n/a

### v1

@@ VI CONCLUSION
Deep-learning vision is advancing rapidly due to performance improvements of neural networks, but its use for the automation of recovery tasks such as waste sorting and disassembly remains minimal. The adaptability and scalability of deep-learning vision can enable highly-flexible waste sorting facilities to process very different types of materials and products. This is currently not feasible with traditional automated methods, which can process only particular classes of items. In addition, deep-learning vision can be used for wide-area autonomous materials monitoring to improve natural resources management (as detailed in [106]).
@@
The term “circular” itself implies a systemic perspective that considers the material life-cycle from extraction to its reuse and disposal. Hence, this review proposes compartmental dynamical thermodynamics to establish rigorous physics-based theoretical foundations of circular flow designs. Systems such as robots, transportation systems, and incinerators, obey the laws of thermodynamics, and hence, their interactions form a network of thermodynamic compartments. The thermodynamic compartments can be added, modified, or removed at the design stage of the compartmental network in order to achieve a circular material flow. As we showed in this review, a theoretically-coherent physics-based framework underpinning the design of circular flow systems is missing in the literature and it is needed to speed-up the implementation of circularity at the local and global scales.
@@ ACKNOWLEDGMENT
F. Zocco sincerely thanks the Department of Information Engineering and Mathematics of University of Siena for hosting him for this research.

### v2

@@ VII CONCLUSION
In this paper, we merged deep-learning vision, compartmental dynamical thermodynamics, and robotic manipulation with circular economy. Deep-learning vision is advancing rapidly due to performance improvements via neural networks, but its use for the automation of recovery tasks such as waste sorting and disassembly remains relatively undeveloped. The adaptability and scalability of deep-learning vision can enable highly-flexible waste sorting facilities to process very different types of materials and products. This is currently not feasible with traditional automated methods, which can process only particular classes of items. In addition, deep-learning vision can be used for wide-area autonomous materials monitoring to improve natural resources management (as detailed in [116]).
@@
The term “circular” itself implies a systemic perspective that considers the material life-cycle from extraction to its reuse and disposal. Hence, this paper proposes compartmental dynamical thermodynamics to establish rigorous physics-based theoretical foundations of circular flow designs. Systems such as robots, transportation systems, and incinerators, obey the laws of thermodynamics, and hence, their interactions form a network of thermodynamic compartments. The thermodynamic compartments can be added, modified, or removed at the design stage of the compartmental network in order to achieve a circular material flow. As we showed in this paper, a theoretically-coherent physics-based framework underpinning the design of circular flow systems is missing in the literature and it is needed to speed-up the implementation of circularity at the local and global scales.
The application of the proposed framework to the case study of plastics systems highlighted several points. Firstly, the framework provides a clear and intuitive representation of supply-recovery chains, especially for those with an industrial and thermal engineering expertise since the framework results from the generalization of the design of thermodynamic cycles such as the Rankine cycle. Secondly, the framework is flexible and scalable, and hence, it enables the consideration of existing network configurations along with simulating hypothetical architectures as needed in what-if analysis. Thirdly, compared to existing frameworks for CE, this framework can coherently integrate robotics and autonomous systems since it is predicated on dynamical systems modeling rather than data analysis as in traditional approaches to CE, e.g., MFA [31]. Fourthly, RL controllers have great potential for making robotic manipulators as flexible as humans, but currently it is an unexplored area in the literature for waste sorting and disassembly operations. As a result, its development requires to work not only on finding and tuning a suitable RL algorithm, but also creating the RL simulation and testing environment for CE applications. Fifthly, the conceptualization of CE as an optimization problem as in Equation (2) highlights that variations occurring within a single compartment need to be propagated to the other compartments until the unsustainable flows are reached in order to quantify the impact of the compartmental variation on the network circularity. This is because there is a strong interrelationship between compartment-level dynamics and the overall network circularity. Sixthly, Equation (2) also has highlighted the similarity and the difference between compartmental and artificial neural networks. While both are designed as a result of an optimization problem, their constituent units are fundamentally different; with compartments being employed in the former and the neurons in the latter.
Future work will consist on developing RL environments and algorithms for CE applications such as robotic waste sorting and disassembly; this could be followed by their experimentation on real manipulators. Another future research direction could be to simulate the plastics systems discussed in the case study in order to analyse existing and hypothetical configurations in collaboration with industry and supply-chain practitioners.
@@ ACKNOWLEDGMENT
F. Zocco sincerely thanks the Department of Information Engineering and Mathematics of University of Siena (Italy) for hosting him for this research and Mohd Shoaib at Nottingham Trent University (UK) for the valuable discussions on reinforcement learning.